if μΔλ= μ+λ/μ-λ’, what is the value of 6/7Δ7/6?
Answer by greenestamps(13200)



Not clear....

μ+λ/μ-λ’ = μ + (λ/μ) - λ’ ?
μ+λ/μ-λ’ = (μ+λ)/(μ-λ’) ?
And I have no idea what the " ’ " means....

If you want help with this, re-post in a way that we know what the question is.